# Galaxy And Mass Assembly (GAMA): from filaments to voids, how extreme environment affects gas metallicity and SFR in galaxies

### v1

@@ 5 Conclusions
This study shows that the environment affects galaxy metallicity and SFR. The main conclusions are as follows.
Galaxies in filaments exhibit higher metallicity and lower star formation rates due to their enriched environments and ongoing interactions.
Void galaxies have the lowest metallicity but show high star formation rates driven by recurrent bursts fueled by halo cold gas reserves and a lack of gas stripping from other galaxies.
The and SFR differences suggest that the local environment, rather than large-scale membership, has a greater influence on galaxies’ evolutionary properties.
@@ Acknowledgements.
We thank the referee for their thoughtful comments that improved the manuscript. This work has been supported by grant PID2021-123417OB-I00 and PCI2022-135023-2 funded by the Spanish Ministry of Science and Innovation. M.A.L.L. acknowledge support from the Ramón y Cajal programme funded by the Spanish Government RYC2020-029354-I. In addition to other packages mentioned, we made direct use of the following Python software packages: astropy (Astropy Collaboration et al. 2013), matplotlib (Hunter 2007), numpy (Harris et al. 2020), and smplotlib (Li 2023); as well as the following R packages: ggplot2 (Wickham 2016), patchwork (Pedersen 2025), and propagate (Spiess 2014). GAMA is a joint European-Australasian project based around a spectroscopic campaign using the Anglo-Australian Telescope. The GAMA input catalogue is based on data from the Sloan Digital Sky Survey and the UKIRT Infrared Deep Sky Survey. Complementary imaging of the GAMA regions is being obtained by a number of independent survey programmes, including GALEX MIS, VST KIDS, VISTA VIKING, WISE, Herschel-ATLAS, GMRT and ASKAP, providing UV to radio coverage. GAMA is funded by the STFC (UK), the ARC (Australia), the AAO and the Participating Institutions. The GAMA website is http://www.gama-survey.org/.References
@@
- Domínguez-Gómez et al. (2022) Domínguez-Gómez, J., Lisenfeld, U., Pérez, I., et al. 2022, A&A, 658, A124
- Domínguez-Gómez et al. (2023a) Domínguez-Gómez, J., Pérez, I., Ruiz-Lara, T., et al. 2023a, A&A, 680, A111
- Domínguez-Gómez et al. (2023b) Domínguez-Gómez, J., Pérez, I., Ruiz-Lara, T., et al. 2023b, Nature, 619, 269
@@
- Torres-Ríos et al. (2024) Torres-Ríos, G., Pérez, I., Verley, S., et al. 2024, in EAS2024, European Astronomical Society Annual Meeting, 1965

### v2

@@ 5 Conclusions
This study shows that the environment affects galaxy metallicity and SFR. The main conclusions are as follows:
Galaxies in filaments exhibit higher metallicity and lower star formation rates, due to their enriched environments and ongoing interactions.
Void galaxies have the lowest metallicity but show high star formation rates driven by recurrent bursts fuelled by halo cold gas reserves and a lack of gas stripping.
The and SFR differences found suggest that the local environment, rather than large-scale membership, has a greater influence on galaxies’ evolutionary properties.
@@ Acknowledgements.
We thank the referee for their thoughtful comments that improved the manuscript. This work has been supported by grant PID2021-123417OB-I00 and PCI2022-135023-2 funded by the Spanish Ministry of Science and Innovation. M.A.L.L. acknowledge support from the Ramón y Cajal programme funded by the Spanish Government RYC2020-029354-I. In addition to other packages mentioned, we made use of: astropy (Astropy Collaboration et al. 2013), matplotlib (Hunter 2007), numpy (Harris et al. 2020), and smplotlib (Li 2023); as well as the following R packages: ggplot2 (Wickham 2016), patchwork (Pedersen 2025), and propagate (Spiess 2014). GAMA is a joint European-Australasian project based around a spectroscopic campaign using the Anglo-Australian Telescope. The GAMA input catalogue is based on data from the Sloan Digital Sky Survey and the UKIRT Infrared Deep Sky Survey. Complementary imaging of the GAMA regions is being obtained by a number of independent survey programmes, including GALEX MIS, VST KIDS, VISTA VIKING, WISE, Herschel-ATLAS, GMRT and ASKAP, providing UV to radio coverage. GAMA is funded by the STFC (UK), the ARC (Australia), the AAO and the Participating Institutions. The GAMA website is http://www.gama-survey.org/.References
@@
- Dom´ınguez-Gómez et al. (2022) Domínguez-Gómez, J., Lisenfeld, U., Pérez, I., et al. 2022, A&A, 658, A124
- Dom´ınguez-Gómez et al. (2023a) Domínguez-Gómez, J., Pérez, I., Ruiz-Lara, T., et al. 2023a, A&A, 680, A111
- Dom´ınguez-Gómez et al. (2023b) Domínguez-Gómez, J., Pérez, I., Ruiz-Lara, T., et al. 2023b, Nature, 619, 269
@@
- Torres-R´ıos et al. (2024) Torres-Ríos, G., Pérez, I., Verley, S., et al. 2024, in EAS2024, European Astronomical Society Annual Meeting, 1965